Is the interaction variation significant in case of the following information concerning mileage based on

different brands of gasoline and cars?


1
Expert's answer
2021-12-20T17:00:12-0500


Two Way ANOVA


"H_0:" there is no interaction between different brands of gasoline and cars

"H_a:" there is an interaction between different brands of gasoline and cars


factor A is cars

factor B is brand of gasoline


F Statistic:

"F_{AB} = MS_{AB} \/ MS_E=2.71"

where:

Mean Square:

MSAB = SSAB / DFAB = 2.71

DFAB = (a - 1)(b - 1) = 6

Sum of Squares:

SSABiaΣjbni,ji,j - Ȳi - Ȳj + Ȳ)2 = 16.25

Error Mean Square:

MSE = SSE / DFE = 1

SSEiaΣjbΣkni,j(Yi,j,k - Ȳi,j)2 = 12


critical value:

"F_{crit}(df_1,df_2)=F_{crit}(6,12)=2.9961"

where error df2:

DFE = n - a*b = "24-3\cdot4=12"


Since "F<F_{crit}" we accept null hypothesis. There is no interaction between different brands of gasoline and cars.